The length of a rectangle is 4 cm longer than its width. If the perimeter of the rectangle is 28 cm, find its area.
Answer by TimothyLamb(4379)

L = W + 4
2L + 2W = 28
---
put the system of linear equations into standard form
---
L - W = 4
2L + 2W = 28
---
replace L with x
replace W with y
---
x - y = 4
2x + 2y = 28
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = L = 9
y = W = 5
---
answer:
a = LW = 9*5 = 45 sq.cm.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php